A patient is to receive an intravenous infusion of 1000 mL of 0.9% Sodium Chloride over 10 hours using a volumetric infusion pump. The IV infusion set has a tubing drop factor of 15 gtt/mL.

The nurse should set the pump at what rate in mL per hour?
Answer #1

Ans) Order: IV Infusion 1000 ml of 0.9% sodium chloride

Time: 10 hours

Drop factor: 15 gtt/ml

Formula:

Rate ( ml/hr) = Volume to be Infused (ml) × drop factor / Time (hr)

= 1000 × 15 / 10 = 15000/10 = 1500 ml/hr.